# A central limit theorem for the stochastic cable equation

### v1

@@ 4 Proof of Theorem 1
$:=\frac{1}{L}1_{(0,t)}(s)1_{[0,L]}(y)\sigma(u(s,y))\int_{0}^{L}G_{t-s}(x,y)\ \mathrm{d}x$
@@
×$\quad\times\mathcal{I}_{0}(t-s_{1},y_{1})\mathcal{I}_{0}(\tau-s_{1},y_{1})\mathcal{I}_{0}(t-s_{2},y_{2})\mathcal{I}_{0}(\tau-s_{2},y_{2})\ \mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2}.$
@@
×$\quad\times\|\sigma(u(s_{2},y_{2}))D_{r,z}\sigma(u(s_{2},y_{2}))\|_{2}\ \mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2}$
@@
×$\quad\times\mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2},$
where $M_{\sigma}$ is the constant in the linear growth condition for $\sigma$. From the chain rule of Malliavin derivative [N06],
$D_{r,z}\sigma(u(s,y))=\sigma^{\prime}(u(s,y))D_{r,z}u(s,y).$
Then, we obtain
@@
×$\quad\times\mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2},$
@@
×$\quad\times\mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2}$
@@
×$\quad\times\mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2},$
@@
×$\quad\times\mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2}$
@@
×$\quad\times\mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2}$
@@
×$\quad\times E[\sigma(u(r_{1},z_{1}))(D_{r_{1},z_{1}}\sigma(u(s,y)))\sigma(u(r_{2},z_{2}))(D_{r_{2},z_{2}}\sigma(u(s,y)))]\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s.$
@@
×$\quad\times\|\sigma(u(r_{2},z_{2}))\|_{4}\|D_{r_{2},z_{2}}\sigma(u(s,y))\|_{4}\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s$
@@
×$\quad\times\|D_{r_{2},z_{2}}\sigma(u(s,y))\|_{4}\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s,$
where $M_{\sigma}$ is the constant in the linear growth condition for $\sigma$. From the chain rule of Malliavin derivative [N06], we obtain
@@
$\leq\frac{e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}}{L^{4}}\int_{0}^{t}\int_{0}^{L}\int_{[0,\tau\wedge s]^{2}}\int_{[0,L]^{2}}\|\sigma^{\prime}(u(s,y))D_{r_{1},z_{1}}u(s,y)\|_{4}$
×$\quad\times\|\sigma^{\prime}(u(s,y))D_{r_{2},z_{2}}u(s,y)\|_{4}\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s$
@@
×$\quad\times\|D_{r_{2},z_{2}}u(s,y)\|_{4}\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s,$
where $K_{\sigma}$ is Lipschitz constant of $\sigma$. In the case of Neumann/Dirichlet boundary conditions, applying Lemma 2.2, we have
@@
×$\quad\times p_{\beta(s-r_{1})}(y-z_{1})p_{\beta(s-r_{2})}(y-z_{2})\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s$
@@
×$\quad\times p_{\beta(s-r_{1})}(y-z_{1})p_{\beta(s-r_{2})}(y-z_{2})\ \mathrm{d}y\mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}s$
@@
×$\quad\times p_{\beta(2s-r_{1}-r_{2})}(z_{1}-z_{2})\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}s,$
@@
×$\quad\times p_{\beta(2s-r_{1}-r_{2})}(z_{1}-z_{2})\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}s$
@@
×$\quad\times G_{s-r_{1}}(y,z_{1})G_{s-r_{2}}(y,z_{2})\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s$
@@
$\sup_{t\in[0,T]}\operatorname{Var}\left(\left\langle DF_{L}(t),v_{L,t}\right\rangle_{\mathcal{H}}\right)\leq\frac{A_{T}}{L^{3}}\text{~{}for~{}all~{}}L\geq 1$
@@
$\leq\frac{2\sqrt{A_{t}}}{L^{3/2}\operatorname{Var}(F_{L}(t))}$
uniformly for all $t\in[0,T]$ and $L\geq 1$. Proposition 3.1 ensures that
$\operatorname{Var}(F_{L}(t))\sim\frac{1}{L}\int_{0}^{t}e^{-2\alpha(t-s)}f_{\sigma}(s)\ \mathrm{d}s$
as $L\to\infty$. This together with (11) implies (4). ∎
@@ 5 Proof of Theorem 2
$\leq 16Le^{2|\alpha|T}\sum_{n=1}^{\infty}\frac{1}{n^{2}\pi^{2}}\int_{0}^{t_{1}}e^{-\frac{n^{2}\pi^{2}s}{L^{2}}}\ \mathrm{d}s\left(|\alpha|^{2}e^{2|\alpha|T}|t_{2}-t_{1}|^{2}+\left|e^{-\frac{n^{2}\pi^{2}(t_{2}-t_{1})}{2L^{2}}}-1\right|^{2}\right)$
$=\frac{16}{6}L|\alpha|^{2}Te^{4|\alpha|T}|t_{2}-t_{1}|^{2}+16Le^{2|\alpha|T}\sum_{n=1}^{\infty}\frac{1}{n^{2}\pi^{2}}\int_{0}^{t_{1}}e^{-\frac{n^{2}\pi^{2}s}{L^{2}}}\ \mathrm{d}s\left|e^{-\frac{n^{2}\pi^{2}(t_{2}-t_{1})}{2L^{2}}}-1\right|^{2}.$
@@
$16Le^{2|\alpha|T}\sum_{n=1}^{\infty}\frac{1}{n^{2}\pi^{2}}\int_{0}^{t_{1}}e^{-\frac{n^{2}\pi^{2}s}{L^{2}}}\ \mathrm{d}s\left|e^{-\frac{n^{2}\pi^{2}(t_{2}-t_{1})}{2L^{2}}}-1\right|^{2}$
$=16Le^{2|\alpha|T}\sum_{n=1}^{\infty}\frac{1}{n^{2}\pi^{2}}\frac{L^{2}}{n^{2}\pi^{2}}\left(1-e^{-\frac{n^{2}\pi^{2}t_{1}}{L^{2}}}\right)\left|e^{-\frac{n^{2}\pi^{2}(t_{2}-t_{1})}{2L^{2}}}-1\right|^{2}$
$\leq 16Le^{2|\alpha|T}\sum_{n=1}^{\infty}\frac{1}{n^{2}\pi^{2}}\left|1\wedge\frac{n^{2}\pi^{2}(t_{2}-t_{1})}{2L^{2}}\right|^{2}\frac{L^{2}}{n^{2}\pi^{2}}$
@@
Therefore, in the case of any boundary conditions, we have (12). This completes the proof. ∎
@@
The tightness of $\{\sqrt{L}F_{L}\}_{L\geq 1}$ in $C([0,T])$ is a consequence of Proposition 5.1 (see [K98]). Then, it remains to prove that the finite-dimensional distributions of the process $(\sqrt{L}F_{L}(t))_{t\in[0,T]}$ converge to those of $(\int_{0}^{t}e^{-\alpha(t-s)}\sqrt{f_{\sigma}(s)}\ \mathrm{d}W_{s})_{t\in[0,T]}$ as $L\to\infty$.
Let us fix $T>0$ and $m\geq 1$ points $t_{1},\dots,t_{m}\in(0,T]$. Proposition 3.1 ensures that, for every $i,j=1,\dots,m$,
@@
as $L\to\infty$. We define the following quantities:
$F_{i}:=\frac{F_{L}(t_{i})}{\sqrt{\operatorname{Var}(F_{L}(t_{i}))}}$
and
$C_{i,j}:=\operatorname{Cov}\left[F_{i},F_{j}\right]$
for $i,j=1,\dots,m$. We will write $F:=(F_{1},\dots,F_{m})$, and let $G=(G_{1},\dots,G_{m})$ denote a centered Gaussian random vector with covariance matrix $C=(C_{i,j})_{1\leq i,j\leq m}$.
Recall from (9) the random fields $v_{L,t_{1}},\dots,v_{L,t_{m}}$, and define rescaled random fields $V_{1},\dots,V_{m}$ as follows:
$V_{i}:=\frac{v_{L,t_{i}}}{\sqrt{\operatorname{Var}(F_{L}(t_{i}))}}$
for $i=1,\dots,m$. According to (8), $F_{i}=\delta(V_{i})$ for all $i=1,\dots,m$. And, by duality, $E[\langle DF_{i},V_{j}\rangle_{\mathcal{H}}]=E[F_{i}\delta(V_{j})]=C_{i,j}$ for all $i,j=1,\dots,m$. Then, Proposition 2.2 implies that
$\left|E[h(F)]-E[h(G)]\right|\leq\frac{1}{2}\|h^{\prime\prime}\|_{\infty}\sqrt{\sum_{i,j=1}^{m}\operatorname{Var}\left(\langle DF_{i},V_{j}\rangle_{\mathcal{H}}\right)}$
for all $h\in C^{2}(\mathbb{R}^{m})$. From Proposition 4.1,
@@
On the other hand, owing to (15), as $L\to\infty$,
$C_{i,j}=\frac{\operatorname{Cov}(F_{L}(t_{i}),F_{L}(t_{j}))}{\sqrt{\operatorname{Var}(F_{L}(t_{i}))}\sqrt{\operatorname{Var}(F_{L}(t_{j}))}}\to\frac{\int_{0}^{t_{i}\wedge t_{j}}e^{-\alpha(t_{i}+t_{j}-2s)}f_{\sigma}(s)\ \mathrm{d}s}{\sqrt{\int_{0}^{t_{i}}e^{-2\alpha(t_{i}-s)}f_{\sigma}(s)\ \mathrm{d}s}\sqrt{\int_{0}^{t_{j}}e^{-2\alpha(t_{j}-s)}f_{\sigma}(s)\ \mathrm{d}s}},$
which yields that the Gaussian random vector $G$ converges weakly to
$\left(\frac{\int_{0}^{t_{1}}e^{-\alpha(t_{1}-s)}\sqrt{f_{\sigma}(s)}\ \mathrm{d}W_{s}}{\sqrt{\int_{0}^{t_{1}}e^{-2\alpha(t_{1}-s)}f_{\sigma}(s)\ \mathrm{d}s}},\dots,\frac{\int_{0}^{t_{m}}e^{-\alpha(t_{m}-s)}\sqrt{f_{\sigma}(s)}\ \mathrm{d}W_{s}}{\sqrt{\int_{0}^{t_{m}}e^{-2\alpha(t_{m}-s)}f_{\sigma}(s)\ \mathrm{d}s}}\right)$ (16)
as $L\to\infty$.
Therefore, it follows that $F$ converges weakly to the random vector in (16) as $L\to\infty$. Then, from (15), it follows that

### v2

@@ 4 Proof of Theorem 1
$=\frac{1}{L}1_{(0,t)}(s)1_{[0,L]}(y)\sigma(u(s,y))\int_{0}^{L}G_{t-s}(x,y)\ \mathrm{d}x$
@@
$\quad\times\mathcal{I}_{0}(t-s_{1},y_{1})\mathcal{I}_{0}(\tau-s_{1},y_{1})\mathcal{I}_{0}(t-s_{2},y_{2})\mathcal{I}_{0}(\tau-s_{2},y_{2})\ \mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2}.$
@@
$\quad\times\|\sigma(u(s_{2},y_{2}))D_{r,z}\sigma(u(s_{2},y_{2}))\|_{2}\ \mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2}$
@@
$\quad\times\mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2},$
where $M_{\sigma}$ is the constant in the linear growth condition for $\sigma$. From the chain rule of Malliavin derivative for a Lipschitz function with constant $K_{\sigma}$ [N06],
$D_{r,z}\sigma(u(s,y))=G_{\sigma,u(s,y)}D_{r,z}u(s,y),$
where $G_{\sigma,u(s,y)}$ is a random variable bounded by $K_{\sigma}$. Then, we obtain
@@
$\quad\times\mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2},$
@@
$\quad\times\mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2}$
@@
$\quad\times\mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2},$
@@
$\quad\times\mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2}$
@@
$\quad\times\mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2}$
@@
$\quad\times E[\sigma(u(r_{1},z_{1}))(D_{r_{1},z_{1}}\sigma(u(s,y)))\sigma(u(r_{2},z_{2}))(D_{r_{2},z_{2}}\sigma(u(s,y)))]\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s.$
@@
$\quad\times\|\sigma(u(r_{2},z_{2}))\|_{4}\|D_{r_{2},z_{2}}\sigma(u(s,y))\|_{4}\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s$
@@
$\quad\times\|D_{r_{2},z_{2}}\sigma(u(s,y))\|_{4}\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s,$
where $M_{\sigma}$ is the constant in the linear growth condition for $\sigma$. From the chain rule of Malliavin derivative for a Lipschitz function [N06], we obtain
@@
$\leq\frac{e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}}{L^{4}}\int_{0}^{t}\int_{0}^{L}\int_{[0,\tau\wedge s]^{2}}\int_{[0,L]^{2}}\|G_{\sigma,u(s,y)}D_{r_{1},z_{1}}u(s,y)\|_{4}$
$\quad\times\|\sigma^{\prime}(u(s,y))D_{r_{2},z_{2}}u(s,y)\|_{4}\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s$
@@
$\quad\times\|D_{r_{2},z_{2}}u(s,y)\|_{4}\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s,$
where $K_{\sigma}$ is Lipschitz constant of $\sigma$ and $G_{\sigma,u(s,y)}$ is a random variable bounded by $K_{\sigma}$. In the case of Neumann/Dirichlet boundary conditions, applying Lemma 2.2, we have
@@
$\quad\times p_{\beta(s-r_{1})}(y-z_{1})p_{\beta(s-r_{2})}(y-z_{2})\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s$
@@
$\quad\times p_{\beta(s-r_{1})}(y-z_{1})p_{\beta(s-r_{2})}(y-z_{2})\ \mathrm{d}y\mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}s$
@@
$\quad\times p_{\beta(2s-r_{1}-r_{2})}(z_{1}-z_{2})\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}s,$
@@
$\quad\times p_{\beta(2s-r_{1}-r_{2})}(z_{1}-z_{2})\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}s$
@@
$\quad\times G_{s-r_{1}}(y,z_{1})G_{s-r_{2}}(y,z_{2})\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s$
@@
$\sup_{t\in[0,T]}\operatorname{Var}\left(\left\langle DF_{L}(t),v_{L,t}\right\rangle_{\mathcal{H}}\right)\leq\frac{A_{T}}{L^{3}}$
@@
$\leq\frac{2\sqrt{A_{t}}}{L^{3/2}\operatorname{Var}(F_{L}(t))}.$
From Proposition 3.1, the asymptotic behavior of the variance is given by
$\operatorname{Var}(F_{L}(t))\gtrsim\frac{1}{L}\frac{\sigma(1)^{2}}{2}\int_{0}^{t\wedge\delta}e^{-2\alpha(t-s)}\ \mathrm{d}s,$
as $L\to\infty$. Consequently, using this result along with $\sigma(1)\neq 0$ and (19), we can deduce (4). ∎
@@ 5 Proof of Theorem 2
[…]
@@
[…]
@@
Therefore, in the case of Neumann, Dirichlet, or periodic boundary conditions, we have (20). This completes the proof. ∎
@@
The tightness of the family of processes $\{\sqrt{L}F_{L}(\cdot)\}_{L\geq 1}$ in $C([0,T])$ is guaranteed by Proposition 5.1 (see, e.g., [K98]). What remains is to show that the finite-dimensional distributions of $(\sqrt{L}F_{L}(t))_{t\in[0,T]}$ converge to those of $(\int_{0}^{t}e^{-\alpha(t-s)}\sqrt{f_{\sigma}(s)}\ \mathrm{d}W_{s})_{t\in[0,T]}$ as $L\to\infty$
Let us fix $T>0$ and $m\geq 1$ points $t_{1},\dots,t_{m}\in(0,T]$. According to Proposition 3.2, as $L\to\infty$, the covariance satisfies
@@
for all $i,j=1,\dots,m$. To proceed, let us define the vector $F:=(F_{1},\dots,F_{m})$ with components
$F_{i}:=\frac{F_{L}(t_{i})}{\sqrt{\operatorname{Var}(F_{L}(t_{i}))}}.$
We then define $G=(G_{1},\dots,G_{m})$ to be a centered Gaussian random vector whose covariance matrix $C=(C_{i,j})$ is given by
$C_{i,j}:=\operatorname{Cov}\left[F_{i},F_{j}\right].$
Let us introduce the rescaled random fields $V_{1},\dots,V_{m}$ by setting
$V_{i}:=\frac{v_{L,t_{i}}}{\sqrt{\operatorname{Var}(F_{L}(t_{i}))}},\quad i=1,\dots,m,$
where the fields $v_{L,t_{i}}$ are given in (17). From (16), we have the relation $F_{i}=\delta(V_{i})$. By duality, $E[\langle DF_{i},V_{j}\rangle_{\mathcal{H}}]=E[F_{i}\delta(V_{j})]=C_{i,j}$ for all $i,j=1,\dots,m$. Applying Proposition 2.2 then yields the desired bound for any $h\in C^{2}(\mathbb{R}^{m})$:
$\left|E[h(F)]-E[h(G)]\right|\leq\frac{1}{2}\|h^{\prime\prime}\|_{\infty}\sqrt{\sum_{i,j=1}^{m}\operatorname{Var}\left(\langle DF_{i},V_{j}\rangle_{\mathcal{H}}\right)}.$
From Proposition 4.1,
@@
We also examine the limit of the covariances $C_{i,j}$. Using the result from (23), we find that as $L\to\infty$,
$C_{i,j}\to\frac{\int_{0}^{t_{i}\wedge t_{j}}e^{-\alpha(t_{i}+t_{j}-2s)}f_{\sigma}(s)\ \mathrm{d}s}{\sqrt{\int_{0}^{t_{i}}e^{-2\alpha(t_{i}-s)}f_{\sigma}(s)\ \mathrm{d}s}\sqrt{\int_{0}^{t_{j}}e^{-2\alpha(t_{j}-s)}f_{\sigma}(s)\ \mathrm{d}s}}.$
It follows that the law of the Gaussian vector $G$, which is determined by its covariance structure, converges weakly to that of
$\left(\frac{\int_{0}^{t_{1}}e^{-\alpha(t_{1}-s)}\sqrt{f_{\sigma}(s)}\ \mathrm{d}W_{s}}{\sqrt{\int_{0}^{t_{1}}e^{-2\alpha(t_{1}-s)}f_{\sigma}(s)\ \mathrm{d}s}},\dots,\frac{\int_{0}^{t_{m}}e^{-\alpha(t_{m}-s)}\sqrt{f_{\sigma}(s)}\ \mathrm{d}W_{s}}{\sqrt{\int_{0}^{t_{m}}e^{-2\alpha(t_{m}-s)}f_{\sigma}(s)\ \mathrm{d}s}}\right).$ (24)
Therefore, $F$ converges weakly to the random vector in (24) as $L\to\infty$. Recalling the definition of $F_{i}$ and using the asymptotic variance from (23), the random vector